# Resonant Photon-Axion Mixing Driven by Dark Matter Oscillations

### v1

@@ IV Numerical analysis
For each parameter pair $(m_{a},g_{a\gamma})$, we solved the equations of motion 1,000 times, using equally spaced values between 0 and $2\pi$ for the initial axion phase $\varphi_{a}$ and assuming a $30\%$ initial linear polarization [Liodakis:2021els], then we obtained a well-resolved ensemble of induced circular polarization. In Fig. S2, the resulting probability density of the degree of circular polarization $\Pi_{C}$ is visualized using the kernel density estimation. By comparing the variance of this predicted signal with characteristic observational uncertainties, we set conservative consistency limits on the axion–photon coupling $g_{a\gamma}$ for the mass range $10^{-23}-10^{-20}\,\mathrm{eV}$ in Fig. 2.

### v2

@@ IV Numerical analysis
For each parameter pair $(m_{a},g_{a\gamma})$, we solved the equations of motion 1,000 times, using equally spaced values between 0 and $2\pi$ for the axion phase $\varphi_{a}$ and assuming a $30\%$ initial linear polarization [26], then we obtained a well-resolved ensemble of induced circular polarization. In Fig. S2, the resulting probability density of the degree of circular polarization $\Pi_{C}$ is visualized using the kernel density estimation.
To derive constraints on the axion–photon coupling $g_{a\gamma}$, we adopt a source-agnostic statistical criterion: the standard deviation of the axion-induced circular polarization $\sigma_{\mathrm{std}}$ must not exceed the characteristic observational uncertainty of a given source. For the case of 3C 279, we take $\sigma_{\Pi_{C}}\approx 0.30\%$ [26]. This criterion ensures consistency between the predicted stochastic signal and the measured noise level without assuming a deterministic phase alignment. For any astrophysical source, the axion phase $\varphi_{a}$ is effectively constant over observational timescales. However, since its value is a priori unknown, the axion-induced circular polarization manifests as a stochastic contribution with vanishing mean. Even when the axion-induced circular polarization exhibits a bimodal distribution for certain masses, its sign and magnitude remain observationally unpredictable due to the unknown axion phase, making the variance a physically meaningful measure of its impact. Requiring its variance not to exceed the instrumental sensitivity therefore provides a consistency bound that is insensitive to detailed phase realizations. Figure 2 shows the resulting illustrative bound in parameter space, with CAST limits [5] included for scale comparison.
For $m_{a}\in[10^{-23},10^{-20}]\,\mathrm{eV}$, this procedure yields limits at the level of $g_{a\gamma}\lesssim 2\times 10^{-11}\,\mathrm{GeV}^{-1}$. The distinctive band-like structure of the resulting contour is a characteristic fingerprint of the Floquet-driven mode mixing mechanism, requiring driven, time-dependent mixing and the associated discrete resonance structure. Two generic features emerge from this mechanism: (1) The conversion efficiency, affected by the Bessel functions $J_{n}(\Delta_{\text{DM}}/m_{a})$, is maximized when the effective momentum mismatch is bridged by integer harmonics of $m_{a}$, a direct consequence of the periodic Floquet driving. The sharp dips in the consistency bound correspond to resonance “sweet spots” where this condition is optimally satisfied, largely independent of the detailed jet profile. (2) As $m_{a}$ varies, the dominant contribution shifts discretely between different resonance orders $n$, producing the characteristic banded pattern. Weaker constraints between bands reflect off-resonance regimes, where the system approaches the adiabatic limit and the driven mixing becomes inefficient, resulting in suppressed circular polarization production.
Although the current limit region is normalized to the specific noise floor of 3C 279, the qualitative structure of the constraint reflects a phenomenological consequence of the underlying mechanism, thereby illustrating the physical relevance of the framework.